OpenStudy (anonymous):

Yvonna put $4,000 in a savings account. At the end of 3 years, the account had earned $960 in simple interest. a. How much does she have in her account by the end of 3 years? b. At what annual simple interest rate did the account grow? Show your work. c. How many more dollars would she have in her account if the interest rate were 1% greater? Show your work.

OpenStudy (anonymous):

\[S.I = \frac{ P \times R \times R \times T }{ 100 }\] S.I=Simple Interest P=Principal R=Rate T=Time